Grade 10Analytical Geometry

A ray of light travels along a line y = 4 and strikes the surface of a curve
y2 = 4x(x + y) then equation of the line along reflected ray travel is

(a) x = 0

(b) x = 2

(c) x + y – 4

(d) 2x + y = 4

Hrishant Goswami
12 Years agoGrade 10

1 Answer

Amit Saxena
12 Years ago

(A)

 

Sol.     Given curse is (y –2)2 = 4(x + 1)

 

            focus (x + 1 = 1, y – 2 = 0)               Þ        (0, 2)

 

            Point of intersection of the curve and

y = 4 is (0, 4) from the reflection property of

parabola reflected ray passes through the focus.

 

            \         x = 0 is required line